OpenStudy (anonymous):

solve the following system of equations \[\large 1+a^3+3ab=b^3 \\ \large 1+a^5=b^5\] a,b are real numbers …

OpenStudy (anonymous):

hint \[\large x^3 + y^3 + z^3 - 3xyz = (x + y + z) (x^2 + y^2 + z^2 - xy -yz - zx)\]

OpenStudy (anonymous):

From first by using this formula, I got: b -a = 1 What to do next??

OpenStudy (anonymous):

put it in second one

OpenStudy (anonymous):

Second equation??? But how, I have no idea..

OpenStudy (anonymous):

b=a+1 1+a^5=(a+1)^5

OpenStudy (anonymous):

Not sure but it becomes: \[(a+1)^5 - a^5 = 1\] Is there any formula then??

OpenStudy (anonymous):

Sorry it can also become as: \[(a + 1)^5 = a^5 + 1^5\]

OpenStudy (anonymous):

use binomial theorem to expand (a+1)^5

OpenStudy (anonymous):

Oh I am also thinking the same..

OpenStudy (anonymous):

\[(a+1)^5= a^{10} + 5a^4 + 10a^3 + 10a^2 +5a +32\] Am I right??

OpenStudy (anonymous):

1 not 32

OpenStudy (anonymous):

and a^5 not a^10

OpenStudy (dominusscholae):

Hint: There will be two values for a. (Is this correct?)

OpenStudy (anonymous):

Oh sorry.. \[(a+1)^5= a^5 + 5a^4 + 10a^3 + 10a^2 +5a +1\] Is this right now??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So we have left with: \[5a^4 + 10a^3 + 10a^2 +5a = 0\]

OpenStudy (anonymous):

Now I have to make factors of this??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I am having problem in making factors.. Can they form by grouping these terms or by putting 1. -1 etc etc??

OpenStudy (anonymous):

(0,1) and (-1,0)

OpenStudy (anonymous):

Yeah, I got one factor at a = -1..

OpenStudy (anonymous):

grouping

OpenStudy (anonymous):

u have \[a^4+2a^3+2a^2+a=a^4+a+2a^3+2a^2\]

OpenStudy (anonymous):

I have found it by Long Division Method..

OpenStudy (anonymous):

\[(a+1)(5a^3 + 5a^2 + 5a) = 5a(a + 1)(a^2 + a + 1)\] So, a= 0, a = -1 are the two solutions.. Finding other two also..

OpenStudy (anonymous):

As the question says a and are real, so other two cannot be possible because they are complex.. \[a = \frac{-1 \pm i \sqrt{3}}{2}\]

OpenStudy (anonymous):

@waterineyes well done

OpenStudy (anonymous):

So, a = 0 or it will be -1 According to it, b = 1 or b = 0 So, When a = 0, b = 1 when a = -1 b = 0 are the solutions..

OpenStudy (anonymous):

No, @mukushla you guided me well... Thanks for that really...

OpenStudy (anonymous):

welcome my friend

OpenStudy (anonymous):

Notice fermat's last theorem. x^n + y^n = z^n have no solution(s) for n>2 and (x,y,z) are integers. Applying to eq(2), one of a or b must be 0. and we'll get (0,1)(-1,0). Check eq(1) LHS will be same as RHS. so solutions are (0,1)(-1,0) Is it possible?